OpenStudy (aripotta):

could you combine a^2b√6a and 3a^3b√6a?

OpenStudy (aripotta):

\[a ^{2}b \sqrt{6a}\]+\[3a ^{3}b \sqrt{6a}\]?

OpenStudy (anonymous):

You could take out the common factor, but you will end up with some factors times the sum of two terms. Not really much of an improvement.\[a^2b \sqrt{6a} + 3a^3b \sqrt{6a}=a^2b \sqrt{6a}(1+3a)\]

OpenStudy (aripotta):

but you couldn't add them, right? because in the first one it's a^2 and in the second one it's a^3?

OpenStudy (anonymous):

The key to adding these kinds of terms is to find out if they are like terms. Since the variable parts are not the same, these two are not really like terms, and can't be combined.

OpenStudy (aripotta):

mmk thanks :) that's what i wanted to know